Let T:R^3--->R^4 be defined by T(x1,x2,x3)=(x1+x2, x2+x3, x1-x3, 2x1+x2-x3). Check T is operator. Find kernel and range of T. Find dimensions of kernel.
1
Expert's answer
2020-03-19T05:32:17-0400

First we prove that T is a linear operator:

"T(\begin{pmatrix}
x_1 \\
x_2 \\
x_3
\end{pmatrix}+
\begin{pmatrix}
y_1 \\
y_2 \\
y_3
\end{pmatrix}) = 
T\begin{pmatrix}
x_1 + y_1 \\
x_2 + y_2 \\
x_3 + y_3
\end{pmatrix} = 
\begin{pmatrix}
x_1 + y_1 + x_2 + y_2\\
x_2 + y_2 + x_3 + y_3\\
x_1 + y_1 - x_3 - y_3 \\
2x_1 + 2y_1 + x_2 + y_2 - x_3 - y_3
\end{pmatrix} = 
\\
= \begin{pmatrix}
x_1 + x_2 \\
x_2 + x_3 \\
x_1 - x_3 \\
2x_1 + x_2 - x_3
\end{pmatrix} + 
\begin{pmatrix}
y_1 + y_2\\
y_2 + y_3\\
y_1 - y_3 \\
2y_1 + y_2 - y_3
\end{pmatrix} = 
T\begin{pmatrix}
x_1 \\
x_2 \\
x_3
\end{pmatrix}+
T\begin{pmatrix}
y_1 \\
y_2 \\
y_3
\end{pmatrix}
\\
T(\alpha\begin{pmatrix}
x_1 \\
x_2 \\
x_3
\end{pmatrix}) = 
T(\begin{pmatrix}
\alpha x_1 \\
\alpha x_2 \\
\alpha x_3
\end{pmatrix}) = 
\begin{pmatrix}
\alpha x_1 + \alpha x_2 \\
\alpha x_2 + \alpha x_3 \\
\alpha x_1 - \alpha x_3 \\
2\alpha x_1 + \alpha x_2 - \alpha x_3
\end{pmatrix} = 
\begin{pmatrix}
\alpha(x_1 + x_2) \\
\alpha(x_2 + x_3) \\
\alpha(x_1 - x_3) \\
\alpha(2x_1 + x_2 - x_3)
\end{pmatrix} = 
\\
= \alpha\begin{pmatrix}
x_1 + x_2 \\
x_2 + x_3 \\
x_1 - x_3 \\
2x_1 + x_2 - x_3
\end{pmatrix} = 
\alpha T\begin{pmatrix}
x_1 \\
x_2 \\
x_3
\end{pmatrix}
\\"

Then we find the kernel:

"\begin{pmatrix}
x_1 + x_2 \\
x_2 + x_3 \\
x_1 - x_3 \\
2x_1 + x_2 - x_3
\end{pmatrix} = 
\begin{pmatrix}
0 \\
0 \\
0 \\
0
\end{pmatrix}
\\
\begin{pmatrix}
x_1 + x_2 \\
x_2 + x_3 \\
x_1 - x_3 \\
2x_1 + x_2 - x_3
\end{pmatrix} = 
\begin{pmatrix}
1 & 1 & 0 \\
0 & 1 & 1 \\
1 & 0 & -1 \\
2 & 1 & -1
\end{pmatrix}
\begin{pmatrix}
x_1 \\
x_2 \\
x_3
\end{pmatrix}
\\
\begin{pmatrix}
1 & 1 & 0 \\
0 & 1 & 1 \\
1 & 0 & -1 \\
2 & 1 & -1
\end{pmatrix} = 
\begin{pmatrix}
1 & 1 & 0 \\
1 & 1 & 0 \\
1 & 0 & -1 \\
1 & 1 & 0
\end{pmatrix} = 
\begin{pmatrix}
1 & 1 & 0 \\
1 & 0 & -1 \\
0 & 0 & 0 \\
0 & 0 & 0
\end{pmatrix} = 
\begin{pmatrix}
1 & 1 & 0 \\
0 & 1 & 1 \\
0 & 0 & 0 \\
0 & 0 & 0
\end{pmatrix}
\\
x_2 = -x_3 \\
x_1 = -x_2 = x_3"

So the kernel is all vectors given as"\begin{pmatrix}
x_3\\
-x_3 \\
x_3
\end{pmatrix}" . Its dimension is 1, its basis is "\begin{pmatrix}
1 \\
-1 \\
1
\end{pmatrix}" .

Finally we find the range of T:

"\begin{pmatrix}
x_1 + x_2 \\
x_2 + x_3 \\
x_1 - x_3 \\
2x_1 + x_2 - x_3
\end{pmatrix} = 
\begin{pmatrix}
x_1 \\
0 \\
x_1 \\
2x_1
\end{pmatrix} + 
\begin{pmatrix}
x_2 \\
x_2 \\
0 \\
x_2
\end{pmatrix} + 
\begin{pmatrix}
0 \\
x_3 \\
-x_3 \\
-x_3
\end{pmatrix}"

Check if the vectors are linearly independent:

"\begin{pmatrix}
1 & 0 & 1 & 2 \\
1 & 1 & 0 & 1 \\
0 & 1 & -1 & -1
\end{pmatrix} = 
\begin{pmatrix}
1 & 0 & 1 & 2 \\
0 & 1 & -1 & -1 \\
0 & 1 & -1 & -1
\end{pmatrix} = 
\begin{pmatrix}
1 & 0 & 1 & 2 \\
0 & 1 & -1 & -1 \\
0 & 0 & 0 & 0
\end{pmatrix}"

So its basis is "\begin{pmatrix}
1 \\
0 \\
1 \\
2
\end{pmatrix}, 
\begin{pmatrix}
0 \\
1 \\
-1 \\
-1
\end{pmatrix}" .